OpenStudy (anonymous):

Anton joined a golf club two years ago. he pays annual fee of $895 and a fee of $30 each time he plays a game of golf. The function below can be used to calculate the totql yearly golfing fee, f(g)=895+30g where g represents the number of times he played golf during the year. Last year he paid $2,065 as a total golfing fee. for how many games did he pay a greens fee? please explain.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Anton joined a golf club two years ago. he pays annual fee of $895 and a fee of $30 each time he plays a game of golf. The function below can be used to calculate the totql yearly golfing fee, f(g)=895+30g where g represents the number of times he played golf during the year. Last year he paid $2,065 as a total golfing fee. for how many games did he pay a greens fee? please explain.

OpenStudy (anonymous):

here we have 2065=895+30g or g=(2065-895)/30=39